# Robust Perception-Informed Navigation using PAC-NMPC with a Learned Value Function

### v2

@@ Abstract
Nonlinear model predictive control (NMPC) is typically restricted to short, finite horizons to limit the computational burden of online optimization. As a result, global planning frameworks are frequently necessary to avoid local minima when using NMPC for navigation in complex environments. By contrast, reinforcement learning (RL) can generate policies that minimize the expected cost over an infinite-horizon and can often avoid local minima, even when operating only on current sensor measurements. However, these learned policies are usually unable to provide performance guarantees (e.g., on collision avoidance), especially when outside of the training distribution. In this paper, we augment Probably Approximately Correct NMPC (PAC-NMPC), a sampling-based stochastic NMPC algorithm capable of providing statistical guarantees of performance and safety, with an approximate perception-dependent value function trained via RL. We demonstrate in simulation that our algorithm can improve the long-term behavior of PAC-NMPC while outperforming other approaches with regards to safety for both planar car dynamics and more complex, high-dimensional fixed-wing aerial vehicle dynamics. We also demonstrate that, even when our value function is trained in simulation, our algorithm can successfully achieve statistically safe navigation on hardware using a 1/10th scale rally car in cluttered real-world environments using only current sensor information.
@@ I Introduction
In this paper, we augment PAC-NMPC [7] with an RL-trained perception-informed value function to achieve perception-based navigation with statistical performance guarantees even when the learned value function is approximate and uncertain. Using Monte Carlo dropout to capture network uncertainty, we learn a stochastic model of the value function. We then use PAC-NMPC to minimize a bound on the expected terminal costs and constraints derived from this value function. This encourages our receding-horizon NMPC algorithm to assume the long-rage behavior of the RL policy, while preserving the statistical safety guarantees afforded by minimizing the sample-complexity bounds derived in [16]. We demonstrate through simulation and hardware experiments that our algorithm can achieve probabilistically safe perception-informed navigation, improved sim-to-real transfer, robustness to out-of-distribution scenarios, and scales to complex, nonlinear high-dimensional dynamic systems.
@@ II Related Work
To the best of our knowledge, our approach is the first to combine a learned perception-dependent stochastic value function with NMPC to achieve probabilistic guarantees for navigation and obstacle avoidance.
@@ III-A PAC-NMPC
More specifically, PAC-NMPC optimizes a time-varying local feedback policy of the form $\mathbf{u}_{t}=\mathbf{K}_{t}(\bm{\tau}^{d}(\bm{\xi},\mathbf{x}_{0}))\left(\mathbf{x}_{t}^{d}(\bm{\xi},\mathbf{x}_{0})-\mathbf{x}_{t}\right)+\mathbf{u}_{t}^{d}(\bm{\xi})$ where $\bm{\tau}^{d}\triangleq\{\mathbf{x}_{0}^{d},\mathbf{u}_{0}^{d},\mathbf{x}_{1}^{d},\mathbf{u}_{1}^{d},...,\mathbf{u}_{N_{T}-1}^{d},\mathbf{x}_{N_{T}}^{d}\}$ is the nominal trajectory computed using a nominal deterministic discrete-time dynamics model $\mathbf{x}_{t+1}^{d}=\mathbf{x}_{t}^{d}+\mathbf{f}(\mathbf{x}_{t}^{d},\mathbf{u}_{t}^{d})\Delta t$ over $N_{T}$ time steps. Here $\mathbf{x}_{t}^{d}\in\mathbb{R}^{N_{x}}$ and $\mathbf{u}_{t}^{d}\in\mathbb{R}^{N_{u}}$ are the nominal states and control inputs at time index $t$ respectively. $\Delta t$ is the discrete-time time step and $\mathbf{K}_{t}(\bm{\tau}^{d}(\bm{\xi},\mathbf{x}_{0}))\in\mathbb{R}^{N_{u}\times N_{x}}$ is a sequence of time-varying feedback gains computed using the finite horizon, discrete, time-varying linear quadratic regulator (TVLQR). The policy is parameterized only by the nominal input sequence, so that $\bm{\xi}~{}=[{\mathbf{u}^{d}_{0}}^{T}\ {\mathbf{u}^{d}_{1}}^{T}\ ...\ {\mathbf{u}^{d}_{N_{T}-1}}^{T}]^{T}$. The surrogate distribution, $p(\bm{\xi}|\bm{\nu})$, is parameterized as a multivariate Gaussian, $\mathcal{N}\left(\bm{\xi}|\bm{\mu},\bm{\Sigma}\right)$, with a mean, $\bm{\mu}$, and a covariance matrix, $\bm{\Sigma}$. Since the covariance matrix is diagonal, the hyperparameters can be written as $\bm{\nu}\triangleq[\bm{\mu}^{T}\ diag(\bm{\Sigma})^{T}]^{T}$.
For a discrete-time trajectory sequence $\bm{\tau}=\{\mathbf{x}_{0},\mathbf{u}_{0},\mathbf{x}_{1},\mathbf{u}_{1}...,\mathbf{u}_{N_{T}-1},\mathbf{x}_{N_{T}}\}$, one can define a non-negative trajectory cost function, $J(\bm{\tau})\geq 0$, and a trajectory constraint violation function $C(\bm{\tau})\in\{0,1\}$. During each planning interval, PAC-NMPC iteratively optimizes $\bm{\nu}^{*}=\operatorname*{arg\,min}_{\bm{\nu}}\min_{\alpha>0}(\mathcal{J}^{+}_{\alpha}(\bm{\nu})+\gamma\mathcal{C}^{+}_{\alpha}(\bm{\nu}))$, where $\mathcal{J}^{+}_{\alpha}(\bm{\nu})$ is the PAC bound on $J(\bm{\tau})$, $\mathcal{C}^{+}_{\alpha}(\bm{\nu})$ is the PAC bound on $C(\bm{\tau})$, and $\gamma>0$ is a heuristically selected weighting coefficient. These PAC bounds take the form
@@ III-B Actor Critic Reinforcement Learning
Actor Critic algorithms are policy gradient approaches to solve Markov decision processes (MDP) by simultaneously learning a policy and action-value function [47]. Consider a MDP with states $\mathbf{s}_{t}\in\mathbb{R}^{N_{s}}$, actions $\mathbf{a}_{t}\in\mathbb{R}^{N_{a}}$, transition dynamics $p(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t})$, and reward function $r(\mathbf{s}_{t},\mathbf{a}_{t})$, where $t$ denotes the time index. The discounted return is defined as the sum of the discounted future rewards of a state-action sequence, $R^{\gamma_{d}}_{t}=\sum_{i=t}^{\infty}\left\{\gamma_{d}^{i-t}r(\mathbf{s}_{i},\mathbf{a}_{i})\right\}$ where, $\gamma_{d}\in[0,1)$ is a discount factor. The action-value function (Q-function), $Q^{\bm{\pi}}(\mathbf{s}_{t},\mathbf{a}_{t})=\mathbb{E}\left[R^{\gamma_{d}}_{t}\middle|\mathbf{s}_{t},\mathbf{a}_{t}\right]$, is the expected discounted return after taking an action $\mathbf{a}_{t}$ from state $\mathbf{s}_{t}$ and then following policy $\mathbf{s}_{t}=\bm{\pi}(\mathbf{a}_{t})$. The value function, $V^{\bm{\pi}}(\mathbf{s}_{t})=Q^{\bm{\pi}}(\mathbf{s}_{t},\bm{\pi}(\mathbf{s}_{t}))$, is the expected discounted return of following a policy, $\mathbf{s}_{t}=\bm{\pi}(\mathbf{a}_{t})$, from a given state $\mathbf{s}_{t}$. Actor-critic algorithms learn parameters $\bm{\phi}$ for an actor policy, $\bm{\pi}^{\bm{\phi}}(\mathbf{s}_{t})$, and parameters $\bm{\psi}$ for a critic action-value function, $Q^{\bm{\psi}}(\mathbf{s}_{t},\mathbf{a}_{t})$, which estimate the optimal policy $\bm{\pi}^{*}(\mathbf{s}_{t})$ and action-value function $Q^{\bm{\pi}^{*}}(\mathbf{s}_{t},\mathbf{a}_{t})$ where $\bm{\pi}^{*}(\mathbf{s}_{t})=\mathbf{a}_{t}^{*}=\operatorname*{arg\,min}_{\mathbf{a}_{t}}\mathbb{E}\left[Q^{\bm{\pi}^{*}}(\mathbf{s}_{t},\mathbf{a}_{t})\right]$ and $Q^{\bm{\pi}^{*}}(\mathbf{s}_{t},\mathbf{a}_{t}^{*})=\mathbb{E}\left[{r(\mathbf{s}_{t},\mathbf{a}^{*}_{t})+\gamma_{d}Q^{\bm{\pi}^{*}}(\mathbf{s}_{t+1},\mathbf{a}_{t+1}^{*})}\right]$.
@@ IV-A Problem Formulation
Here $\mathbf{x}_{l}$ is the lower state bound, $\mathbf{x}_{u}$ is the upper state bound, $r$ is the maximum robot radius. $\mathbf{p}_{o^{j}}$ is an observed occupied point in the world frame, calculated as
@@ V-A Experimental Setup
Our approach considers not only dynamics uncertainty, but also uncertainty in the learned actor and critic networks themselves. For every trajectory sampled by PAC-NMPC, we also sample random dropout masks for the actor and critic networks [53] before evaluating the cost-to-go at the trajectory’s terminal state. This allows us to incorporate uncertainty from value function into the PAC bound computation to provide more accurate performance guarantees.
@@
When running PAC-NMPC, we optimize a 12 timestep trajectory with $\Delta t=0.1$ sec at a replanning period of $H=0.2$ sec, and we interpolate the feedback policies to 50Hz. We set PAC-NMPC parameters as $L=5$, $M=1024$, $\delta=0.05$, and normalize the sampled trajectory costs before optimizing the PAC bounds to achieve tighter bounds. We ran PAC-NMPC with $\gamma=2$ in the cluttered environments and $\gamma=4$ in the concave trap environments. When running MPPI, we similarly normalized the sampled trajectory costs and summed them with the constraints using the same $\gamma$ values. We used the same number of timesteps, $\Delta t$, $H$, number of sampled trajectories, a temperature of $\gamma_{t}=0.35$, and a sampling variance of $\Sigma_{\epsilon}=0.01$. We allowed MPPI to run for as many iterations as possible in the replanning period and interpolated the resulting trajectories to 50Hz.
@@
Figure 3: Simulated cluttered environments. Testing environment where PAC-NMPC fails with only a quadratic terminal cost.
@@ V-D Fixed-wing UAV
To demonstrate that a value function trained on simple dynamics can provide utility for controlling more complex systems, we ran experiments utilizing the bicycle value function to control the fixed-wing. We trained the bicycle value function in these environments with $v_{min}=2$, $v_{max}=8$, and no Gaussian noise. Then, we mapped the fixed-wing state to the bicycle state by extracting $p_{x}$, $p_{y}$, $\theta$, $v$ from $\mathbf{r}$, $\mathbf{q}$, $\mathbf{v}$ and setting $\delta=\operatorname{atan}(\frac{L}{v}\dot{\theta})$. For the terminal cost, we applied the bicycle value function plus $q(\mathbf{x},\mathbf{u})$ where $Q_{\mathbf{r}}=diag([0.00,0.00,0.1])$, $Q_{\bm{\omega}}=diag([0.1,0.1,0.1])$ and $Q_{\bm{\delta}}=diag([1,1,1])$. This outperformed all other approaches, including PAC-NMPC with the fixed-wing value function (Fig. 8). We hypothesize that this may be because the bicycle value function is easier to learn due to its smaller state space.
To demonstrate that our approach is more resilient when the dynamics uncertainty does not match the training environment, we shifted the mean of the Gaussian noise over the body accelerations to $[3.10,2.10,2.11]\frac{m}{s^{2}}$ and $[4.18,2.84,4.71]\frac{rad}{s^{2}}$ and repeated the experiments. Our approach was able to maintain higher success rates with both the fixed-wing value function and the bicycle value function than PAC-NMPC with a quadratic terminal cost or the actor policy (Fig. 9).
@@ VII Discussion & Limitations
Our approach included several limiting assumptions. It assumed the ability to adequately predict future sensor measurements at the terminal NMPC state as well as an accurate stochastic dynamics model. Additionally, while Monte Carlo dropout was able to quantify epistemic model uncertainty, it would be unable to quantify aleatoric uncertainty (i.e., from sensor noise). Future work could explore the application of our approach to more complex environments and dynamics models, the incorporation of perception uncertainty, as well as online refinement of the learned value function.
Acknowledgment
We gratefully acknowledge the support of the Army Research Laboratory under grant W911NF-22-2-0241.

### v3

# I Introduction

@@ Abstract
Nonlinear model predictive control (NMPC) is typically restricted to short, finite horizons to limit the computational burden of online optimization. As a result, global planning frameworks are frequently necessary to avoid local minima when using NMPC for navigation in complex environments. By contrast, reinforcement learning (RL) can generate policies that minimize the expected cost over an infinite-horizon and can often avoid local minima, even when operating only on current sensor measurements. However, these learned policies are usually unable to provide performance guarantees (e.g., on collision avoidance), especially when outside of the training distribution. In this paper, we augment Probably Approximately Correct NMPC (PAC-NMPC), a sampling-based stochastic NMPC algorithm capable of providing statistical guarantees of performance and safety, with an approximate perception-based value function trained via RL. We demonstrate in simulation that our algorithm can improve the long-term behavior of PAC-NMPC while outperforming other approaches with regards to safety for both planar car dynamics and more complex, high-dimensional fixed-wing aerial vehicle dynamics. We also demonstrate that, even when our value function is trained in simulation, our algorithm can successfully achieve statistically safe navigation on hardware using a 1/10th scale rally car in cluttered real-world environments using only current sensor information.
@@ I Introduction
In this paper, we augment Probably Approximately Correct NMPC (PAC-NMPC) [7] with an RL-trained perception-based value function to achieve navigation with statistical performance guarantees even when the learned value function is approximate and uncertain. Using Monte Carlo dropout to capture network uncertainty, we learn a stochastic model of the value function. We then use PAC-NMPC to minimize a bound on the expected terminal costs and constraints derived from this value function. This encourages our receding-horizon NMPC algorithm to assume the long-range behavior of the RL policy, while preserving the statistical safety guarantees afforded by minimizing the sample-complexity bounds derived in [16]. We demonstrate through simulation and hardware experiments that our algorithm can achieve probabilistically safe perception-based navigation, improved sim-to-real transfer, robustness to out-of-distribution scenarios, and scales to complex, nonlinear high-dimensional dynamic systems.
@@ II Related Work
To the best of our knowledge, our approach is the first to combine a learned perception-based stochastic value function with NMPC to achieve probabilistic guarantees for navigation and obstacle avoidance.
@@ III-A PAC-NMPC
More specifically, PAC-NMPC optimizes a time-varying local feedback policy of the form $\mathbf{u}_{t}=\mathbf{K}_{t}(\bm{\tau}^{d}(\bm{\xi},\mathbf{x}_{0}))\left(\mathbf{x}_{t}^{d}(\bm{\xi},\mathbf{x}_{0})-\mathbf{x}_{t}\right)+\mathbf{u}_{t}^{d}(\bm{\xi})$ where $\bm{\tau}^{d}\triangleq\{\mathbf{x}_{0}^{d},\mathbf{u}_{0}^{d},\mathbf{x}_{1}^{d},\mathbf{u}_{1}^{d},...,\mathbf{u}_{N_{T}-1}^{d},\mathbf{x}_{N_{T}}^{d}\}$ is the nominal trajectory computed using a nominal deterministic discrete-time dynamics model $\mathbf{x}_{t+1}^{d}=\mathbf{x}_{t}^{d}+\mathbf{f}(\mathbf{x}_{t}^{d},\mathbf{u}_{t}^{d})\Delta t$ over $N_{T}$ time steps. Here $\mathbf{x}_{t}^{d}\in\mathbb{R}^{N_{x}}$ and $\mathbf{u}_{t}^{d}\in\mathbb{R}^{N_{u}}$ are the nominal states and control inputs at time index $t$ respectively. $\Delta t$ is the discrete time step and $\mathbf{K}_{t}(\bm{\tau}^{d}(\bm{\xi},\mathbf{x}_{0}))\in\mathbb{R}^{N_{u}\times N_{x}}$ is a sequence of time-varying feedback gains computed using the finite horizon, discrete, time-varying linear quadratic regulator (TVLQR). The policy is parameterized only by the nominal input sequence, so that $\bm{\xi}~{}=[{\mathbf{u}^{d}_{0}}^{T}\ {\mathbf{u}^{d}_{1}}^{T}\ ...\ {\mathbf{u}^{d}_{N_{T}-1}}^{T}]^{T}$. The surrogate distribution, $p(\bm{\xi}|\bm{\nu})$, is parameterized as a multivariate Gaussian, $\mathcal{N}\left(\bm{\xi}|\bm{\mu},\bm{\Sigma}\right)$, with a mean, $\bm{\mu}$, and a covariance matrix, $\bm{\Sigma}$. Since the covariance matrix is diagonal, the hyperparameters can be written as $\bm{\nu}\triangleq[\bm{\mu}^{T}\ diag(\bm{\Sigma})^{T}]^{T}$.
For a discrete-time trajectory sequence $\bm{\tau}=\{\mathbf{x}_{0},\mathbf{u}_{0},\mathbf{x}_{1},\mathbf{u}_{1}...,\mathbf{u}_{N_{T}-1},\mathbf{x}_{N_{T}}\}$, one can define a non-negative trajectory cost function, $J(\bm{\tau})\geq 0$, and a trajectory constraint violation function $C(\bm{\tau})\in\{0,1\}$. During each planning interval, PAC-NMPC iteratively optimizes $\bm{\nu}^{*}=\operatorname*{arg\,min}_{\bm{\nu}}\min_{\alpha>0}(\mathcal{J}^{+}_{\alpha}(\bm{\nu})+\gamma\mathcal{C}^{+}_{\alpha}(\bm{\nu}))$, where $\mathcal{J}^{+}_{\alpha}(\bm{\nu})$ is the PAC bound on $J(\bm{\tau})$, $\mathcal{C}^{+}_{\alpha}(\bm{\nu})$ is the PAC bound on $C(\bm{\tau})$, and $\gamma>0$ is a heuristically selected weighting coefficient. $\mathcal{J}^{+}_{\alpha}(\bm{\nu})$ and $\mathcal{C}^{+}_{\alpha}(\bm{\nu})$ take the form
@@ III-B Actor Critic Reinforcement Learning
Actor Critic algorithms are policy gradient approaches to solve Markov decision processes (MDP) by simultaneously learning a policy and action-value function [47]. Consider a MDP with states $\mathbf{s}_{t}\in\mathbb{R}^{N_{s}}$, actions $\mathbf{a}_{t}\in\mathbb{R}^{N_{a}}$, transition dynamics $p(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t})$, and reward function $r(\mathbf{s}_{t},\mathbf{a}_{t})$, where $t$ denotes the time index. The discounted return is defined as the sum of the discounted future rewards of a state-action sequence, $R^{\gamma_{d}}_{t}=\sum_{i=t}^{\infty}\left\{\gamma_{d}^{i-t}r(\mathbf{s}_{i},\mathbf{a}_{i})\right\}$ where, $\gamma_{d}\in[0,1)$ is a discount factor. The action-value function (Q-function), $Q^{\bm{\pi}}(\mathbf{s}_{t},\mathbf{a}_{t})=\mathbb{E}\left[R^{\gamma_{d}}_{t}\middle|\mathbf{s}_{t},\mathbf{a}_{t}\right]$, is the expected discounted return after taking an action $\mathbf{a}_{t}$ from state $\mathbf{s}_{t}$ and then following policy $\mathbf{a}_{t}=\bm{\pi}(\mathbf{s}_{t})$. The value function, $V^{\bm{\pi}}(\mathbf{s}_{t})=Q^{\bm{\pi}}(\mathbf{s}_{t},\bm{\pi}(\mathbf{s}_{t}))$, is the expected discounted return of following a policy, $\mathbf{a}_{t}=\bm{\pi}(\mathbf{s}_{t})$, from a given state $\mathbf{s}_{t}$. Actor-critic algorithms learn parameters $\bm{\phi}$ for an actor policy, $\bm{\pi}^{\bm{\phi}}(\mathbf{s}_{t})$, and parameters $\bm{\psi}$ for a critic action-value function, $Q^{\bm{\psi}}(\mathbf{s}_{t},\mathbf{a}_{t})$, which estimate the optimal policy $\bm{\pi}^{*}(\mathbf{s}_{t})$ and action-value function $Q^{\bm{\pi}^{*}}(\mathbf{s}_{t},\mathbf{a}_{t})$ where $\bm{\pi}^{*}(\mathbf{s}_{t})=\mathbf{a}_{t}^{*}=\operatorname*{arg\,min}_{\mathbf{a}_{t}}\mathbb{E}\left[Q^{\bm{\pi}^{*}}(\mathbf{s}_{t},\mathbf{a}_{t})\right]$ and $Q^{\bm{\pi}^{*}}(\mathbf{s}_{t},\mathbf{a}_{t}^{*})=\mathbb{E}\left[{r(\mathbf{s}_{t},\mathbf{a}^{*}_{t})+\gamma_{d}Q^{\bm{\pi}^{*}}(\mathbf{s}_{t+1},\mathbf{a}_{t+1}^{*})}\right]$.
@@ IV-A Problem Formulation
Here, $\lor$ and $\neg$ are the logical or and not operators, respectively. $\mathbf{x}_{l}$ is the lower state bound, $\mathbf{x}_{u}$ is the upper state bound, $r$ is the maximum robot radius. $\mathbf{p}_{o^{j}}$ is an observed occupied point in the world frame, calculated as
@@ V-A Experimental Setup
Our approach considers not only uncertainty in the dynamics, but also in the learned actor and critic networks. For every trajectory sampled by PAC-NMPC, we also sample dropout masks for the actor and critic networks [53] before evaluating the cost-to-go at the trajectory’s terminal state. Since our approach already requires multiple forward passes through the networks, one for each trajectory, Monte Carlo dropout does not add additional inference time. This allows us to incorporate value function uncertainty into the PAC bound to provide more accurate performance guarantees.
@@
When running PAC-NMPC, we optimize a 12 timestep trajectory with $\Delta t=0.1$ sec at a replanning period of $H=0.2$ sec, and we interpolate the feedback policies to 50Hz. We set PAC-NMPC parameters as $L=5$, $M=1024$, $\delta=0.05$, and normalize the trajectory costs before optimization to achieve tighter PAC bounds. We evaluated PAC-NMPC with several $\gamma$ values and selected the best value for each environment; $\gamma=2$ in the cluttered environments and $\gamma=4$ in the concave trap environments. When running MPPI, we similarly normalized the sampled trajectory costs and summed them with the constraints using the same $\gamma$ values. We used the same number of timesteps, $\Delta t$, $H$, number of sampled trajectories, a temperature of $\gamma_{t}=0.35$, and a sampling variance of $\Sigma_{\epsilon}=0.01$. We allowed MPPI to run for as many iterations as possible in the replanning period and interpolated the resulting trajectories to 50Hz.
@@
Figure 3: Simulated cluttered environments. Testing environment where PAC-NMPC fails when only using a quadratic terminal cost.
@@ V-D Fixed-wing UAV
To demonstrate that a value function trained on simple dynamics can aid in controlling more complex systems, we ran experiments utilizing the bicycle value function to control the fixed-wing. We trained the bicycle value function in these environments with $v_{min}=2$, $v_{max}=8$, and no Gaussian noise. Then, we mapped the fixed-wing state to the bicycle state by extracting $p_{x}$, $p_{y}$, $\theta$, $v$ from $\mathbf{r}$, $\mathbf{q}$, $\mathbf{v}$ and setting $\delta=\operatorname{atan}(\frac{L}{v}\dot{\theta})$. For the terminal cost, we applied the bicycle value function plus $q(\mathbf{x},\mathbf{u})$ where $Q_{\mathbf{r}}=diag([0.00,0.00,0.1])$, $Q_{\bm{\omega}}=diag([0.1,0.1,0.1])$ and $Q_{\bm{\delta}}=diag([1,1,1])$. This outperformed all other approaches, including PAC-NMPC with the fixed-wing value function (Fig. 8). We hypothesize that this may be because the bicycle value function is easier to learn due to its smaller state space.
To demonstrate that our approach is more resilient when the dynamics uncertainty does not match the training environment, we shifted the mean of the noise over the body accelerations to $[3.10,2.10,2.11]\frac{m}{s^{2}}$ and $[4.18,2.84,4.71]\frac{rad}{s^{2}}$ and repeated the experiments. Our approach was able to maintain higher success rates with both the fixed-wing value function and the bicycle value function than PAC-NMPC with a quadratic terminal cost or the actor policy (Fig. 9).
@@ VII Discussion & Conclusion
Our approach included several limiting assumptions. It assumed the ability to adequately predict future sensor measurements at the terminal NMPC state as well as an accurate stochastic dynamics model. Additionally, while Monte Carlo dropout was able to quantify epistemic model uncertainty, it would be unable to quantify aleatoric uncertainty (i.e., from sensor noise). Our approach also can be affected by uncalibrated uncertainty estimates. Future work could explore the incorporation of perception uncertainty, as well as online refinement of the learned value function.